Sodium hydroxide interacted with 273.6g of aluminum sulfate.

Required: mass of precipitate obtained


1
Expert's answer
2021-04-13T02:07:37-0400

Balanced equation

Al2(SO4)3+NaOH"\implies"2 Al(OH)3+3Na2SO4

Given mass of Sodium hydroxide as 273.6g then:

Moles of Al2(SO4)3="\frac{Mass}{MolarMass}"


="\frac{273.6g}{342.15g/moles}"


=0.799"Moles"

Mole ratio of Al2(SO4)3:2Al(OH3)

Therefore Mole Ratio is 1:2


If 1=0.799moles

"\therefore" 2=?


"\frac{2×0.799moles}{1}"


="1.598moles"

But moles of the precipitae are 1.598moles

Therefore Mass=Molar mass×Moles

="78g\/mol\u00d71.598moles"


=124.65g

Mass of precipitate obtained is therefore 124.65g